OpenStudy (anonymous):

In a colony, each of the 20% families have two TV sets, 30% of the rest families have three TV sets. Each of the remaining families has one TV set. Which of the following statement is true’?

hartnn (hartnn):

where are the statements ?

OpenStudy (anonymous):

we would need the answer statements since it is a true or false question

OpenStudy (anonymous):

lets suppose total no of families = x 20% has two sets i.e : 0.2x = 2sets remaining 80% i.e. 0.8x 30% of remaining i.e 0.3*0.8x = 0.24x = 3sets rest i.e x - (0.2x+0.24x) = x - 0.44x = 0.56x = 1set

OpenStudy (anonymous):

(A) 80% families have at least one TV set (B) 50% families have two TV sets (C) 56% families have one TV set (D) 30% families have three TV sets

OpenStudy (anonymous):

in my view (A) and (C) both are true ... what shall be the correct answer ???

hartnn (hartnn):

atleast one means , either 1, or 2 or 3 Each of the remaining families has one TV set ---> 100% families have at least one TV set so, A is FALSE. only C is correct.

OpenStudy (mathmate):



OpenStudy (mathmate):